OpenStudy (emmyma):

Hi! I need some help with a few algebra questions. If anyone could help me, that would be awesome! I'll give you a medal and become a fan. :) - If y varies directly with x and y = -7 when x = -1, find x when y = -84. - Write the point-slope equation of the line through (-2, 1) with a slope of 3. - Write the equation y - 2 = 4(x + 5) in standard form. - Thank you so much! :)

OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. When x=-1 y=-7 -84+77=-7 We changed y with -77 and as it varies directly, x changes with this same amount. -1-77=-78=x Write the point-slope equation of the line through (-2, 1) with a slope of 3. Check the formula, this should not be hard to do... Write the equation y - 2 = 4(x + 5) in standard form. Standard form is y=ax+b Expand the bracket and add 2 to both sides and you have it

OpenStudy (emmyma):

@Andras Thank you so much for your help :)